OpenStudy (anonymous):

First best answer wins a Medal!!!!!!! Note: Remember to show all of the steps that you use to solve the problem. You borrow $500 from your parents to buy a new computer. Your parents charge you 3% simple interest and give you two years to repay the money. a) How much interest do you pay on the loan? b) What is the total amount of money that you owe your parents?

OpenStudy (anonymous):

As i=P*r*t i=500*3\100*2=30 so the 1st ans is $30 and the total amount is 30+500=530

OpenStudy (anonymous):

Use the formula I = Prt, then plug in your values. I = $500 * 3% * 2 I = 500 * 0.03 * 2 Multiply.

OpenStudy (anonymous):

When you multiply the given values you will then get the interest.

OpenStudy (anonymous):

To get the total amount, you will then add 500 to the interest. :) Hope this helps.

OpenStudy (anonymous):

Thanks!!!!!!!